OpenStudy (anonymous):

4/3(3.14)(3^3) v=113.049 cu ft

OpenStudy (ryan123345):

Is that how it looks?

OpenStudy (ipwnbunnies):

I'm pretty sure this is the equation of a the volume of a sphere so it's (4/3)(pi)(3^3)

OpenStudy (ryan123345):

ok i see thanks

OpenStudy (ipwnbunnies):

If you're not supposed to use a calculator, What I can see is cancelling out a 3, then multiply the whole numbers.

OpenStudy (ryan123345):

\[\frac{ 4 }{ 3 }(\pi)(27)\]

OpenStudy (ipwnbunnies):

Right, then a 3 cancels, left with 4(9)pi

OpenStudy (anonymous):

it is the volume of a sphere. I don't know if I solved it the right way. I haven't worked this type of problem before. This is how I got to the solution is v=4/3(3.14)(3^3) 4*3.14/3*27 12.56/3 4.187*27 v=113.049 I have no idea if this is how to solve it or not. I need help to see if I am doing this correctly.

OpenStudy (ipwnbunnies):

You're doing too much in your simplification. You have the numbers laid out right. I suggest to first do 4/3, then multiply everything else.

OpenStudy (anonymous):

Thank you. I have finally finished and passed that class!! It sure was a struggle.